OpenStudy (anonymous):

prove that cos2x=cos^2x-sin^2x

OpenStudy (solomonzelman):

Sure. We know that, (this is the rule) \(\normalsize\color{blue}{ \cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)}\) So, when the a and b are same, (lets say that both "a" and "b" --- are "x" \(\normalsize\color{blue}{ \cos(x+x)=\cos(x)\cos(x)-\sin(x)\sin(x)}\) that is the same as, \(\normalsize\color{blue}{ \cos(x+x)=\cos^2(x)-\sin^2(x)}\)

OpenStudy (anonymous):

No, I won't type now.. :P

OpenStudy (mathmath333):

\(\large\tt \color{black}{\text{u can use formula}}\) \(\large\tt \color{blue}{cos(A+B)=cosAcosB-sinAsinB}\)